A shell and tube heat exchanger with one shell pass and one tube pass will be used to condense the steam to saturated liquid, which enters the shell side as a saturated vapor at 400 K. The tube side contains R-134a refrigerant with an inlet temperature of 300 K and a mean velocity of 0.4 m/s. The steam flow rate is 1.5 kg/s. The tubes are made from AISI 302 stainless steel and have a 1" nominal diameter (Di = 26.6 mm and Do = 33.4 mm). The heat transfer coefficient at the outer surface of the tubes is ho = 10,000 W/m^2K. Assume a fouling resistance of 0.0002 m^2K/W at the inside surface of the tubes and negligible fouling at the outside.

a) If the coolant outlet temperature is not to exceed 320 K, what is the minimum allowable coolant flowrate?

b) How many tubes are required to achieve the desired performance?

c) What tube length is required?
Answer #1

r o They One sheld one tube pass pass - I - to > THA UTIZOS satunded rapon = 4ook shell side Tube side R-1349 > Tci = 300k Tc0.4 K for Turbulent flow inside Tube: Nu = 0.023 Red piz í oftar Beelter Nuz hidi 20.023 (714678331,4% -ezedsion) ln=3 cco1) a= mcpAT Qzi hay 3) 1.5x2138 2 9 X 1485*C 3280300) 6922183 kJ z) M = 107.augls can) of 4ook 1. * Reborigeran How Log mean